A Complete Bibliography of ACM Transactions on Social Computing (TSC)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 April 2024 Version 1.07

Title word cross-reference

19 [LRSW23].

2018 [RFM^{+}18]. 2019 [DK20].

Abuse [TC20]. Abusive [BYSB21]. Accomplishment [FAC20]. Accountable [SAM23]. Accounts [PKK20]. ACM [Cro18]. across [PKK20]. Act [ZYM^{+}21]. Action [MBWM21]. Activism [MADK22]. Activity [SWK^{+}19]. Adoption [MGS20]. Adult [ARZ^{+}22]. Affordance [ZSF22]. Affordances [EC22]. Aggressive [BYSB21]. Airbnb [RCRCOC^{+}19]. Algorithm [LBZZ20]. Alignment [MZ21]. along [HMP^{+}21]. Analysis [CSKD21, HR23, KZB^{+}20, KMV^{+}19, TPR^{+}18]. Analyzing [JWZW18, RMG20]. Anger [BCJ20]. Answer [SWK^{+}19]. API [MYM23]. App [FEH20]. Application [HMP^{+}21]. Applying [ARZ^{+}22]. Approach [Bay18, JM20, TR18]. Archetypes [SWK^{+}19]. Architecture [GP18]. Asia [AC21]. Assessing [YNMA20, SdMAP20]. Assist [FBJF20]. Attention [ZWSS23]. Attributes [MYM23]. Augmented [JCS^{+}21, WFL^{+}19]. Automated [TR18]. Automatic [BYSB21]. Aware [DCDCM19].

Based [KZB^{+}20, PLK^{+}21, CRO^{+}19]. Be [FEH20]. Bedside [SAM23]. Behavior [KPKL23]. Behavioral [ARZ^{+}22]. Bench [SAM23]. Bench-to-Bedside [SAM23].

Benefits [SDCC19]. Between [KPKL23, MZ21]. Beyond [FS19]. Bias [MWY^{+}18]. Biases [BYSB21]. Big [LBZZ20]. Biofeedback [JCS^{+}21]. Blockchain [JB18]. Brazilian [SdMAP20]. Breakdowns [PMY21]. Bringing [KGK^{+}20]. Building [ZYM^{+}21].

Campaigns [Bay18]. Can [FLZ^{+}20, FEH20]. Capital [AC21]. Casa [RCRCOC^{+}19]. Case [FEH20]. centered [MYM23]. Cessation [DMKH23]. Challenges [ARZ^{+}22, KMZ^{+}21]. Channels [KZB^{+}20]. Characteristic [YSBL18]. Cities [SRCGP18]. Citizen [SRP^{+}18, WFL^{+}19]. Civic [SDCC19]. Co [EC22]. Co-opted [EC22]. Collaborative [HR23]. Collective [PKK20, TFLK18]. Comments [KZB^{+}20, PMY21]. Commercial [Bay18]. Communication [SdMAP20, ZSF22]. Communities [TFLK18]. Community [AC21, DMKH23, GP18, KGTA18, LBZZ20, WGB20]. Competition [FKW23]. Computing [Cro18]. Congress [CSKD21]. Connectedness [FBJF20]. Connective [MBWM21]. ConsensUs [LXB^{+}18]. Consequences [KMV^{+}19]. Constrained [MGS20]. Contagion [ARZ^{+}22]. Content [ARN20, MZ21]. Contested [PKK20]. Contests [FS19]. Context [KGK^{+}20]. Contextual [TR18]. Contribution [YNMA20]. Controlled [EC22, KZB^{+}20]. Controversy [GDGM18]. COVID [LRSW23]. COVID-19 [LRSW23]. Crisis [ZWSS23]. Criteria [LXB^{+}18]. Crowd [DCDCM19]. Crowd-Powered [DCDCM19]. Crowdfunding [JM20]. Crowdsourcing [SL18]. Crowdworker [PBL^{+}19]. Crowdworkers [PBL^{+}19]. Cultural [KPKL23]. Cyberbully [CDK18]. Cyberbullying [TR18].

Danmaku [WSH19]. Data [HR23, KMV^{+}19, SAM23]. Data-Sharing [SAM23]. Deadline [DCDCM19]. Deadline-Aware [DCDCM19]. Debate [MADK22]. Decisions [LXB^{+}18]. Declared [MZ21]. Decomposable [SL18]. Deliberative [PMY21]. Design [CG20, FEH20, HMP^{+}21]. Detection [LBZZ20, MYM23, TR18]. Developing [RCRCOC^{+}19, SRCGP18]. Development [MBWM21]. Digital [DMKH23]. Diligence [JM20]. Dimensions [ARN20, JB18]. Disagreement [LXB^{+}18]. Disastrous [FEH20]. Disclosure [LRSW23]. Discourse [EC22]. Discovering [ZYM^{+}21]. Discovery [TH19]. Discussion [TH19]. Disentangle [WBUAV20]. Do [FLZ^{+}20, SdMAP20]. Driven [JWZW18]. Due [JM20]. Duoethnographic [CG20]. During [LRSW23, ZWSS23, MBWM21]. Dyadic [JCS^{+}21]. Dynamics [FKW23, ZWSS23].

Each [DMKH23]. Economy [RCRCOC^{+}19]. Ecosystem [FKW23]. Editorial [RFM^{+}18]. Effective [MSMS19]. Effects [TFLK18]. Elderly [FBJF20]. Emergent [ALL^{+}20]. Emoji [AZ20, FLZ^{+}20, KGK^{+}20, RMG20]. Emote [KZB^{+}20]. Emote-Based [KZB^{+}20]. Emote-Controlled [KZB^{+}20]. Emotions [WBUAV20]. Engagement [DMKH23, LRSW23, SDCC19]. Engineering [HR23]. Entangled [JB18]. Environment [EC22]. Environments [KPKL23]. Ethics [Bay18]. Evaluation [MSMS19, MYM23]. Evaluations [HMP^{+}21]. Events [PKK20, ZWSS23]. Evidence [PMY21]. Evolution [ALL^{+}20]. Examination [PLK^{+}21]. Examining [ARZ^{+}22, ALL^{+}20, RCRCOC^{+}19]. Exchange [RCRCOC^{+}19, SWK^{+}19]. Experience [ARN20, KGTA18]. Experiences [AC21, GB20]. Experimental [GLL^{+}21]. Experiments [FS19]. Expert [CDK18, SL18]. Experts [YNMA20]. Exploring [AZ20, KPKL23, ZWSS23].

Facebook [TC20, WBUAV20, ZYM^{+}21]. Facial [AZ20]. Factors [GLL^{+}21]. Fair [DCDCM19]. False [CRO^{+}19]. Fast [LBZZ20]. Fear [BCJ20, WBUAV20]. Feedback [KZB^{+}20, SLZ^{+}18, TFLK18]. Felt [WBUAV20]. Fitness [CG20]. Formation [PLK^{+}21]. FOSS [GP18]. FOX [LBZZ20]. Framing [MWY^{+}18]. Friend [TC20]. Funder [JM20].

Game [AC21, SRP^{+}18]. Games [KPKL23]. Gender [MZ21]. Gendered [CG20]. General [KMV^{+}19]. generated [ARN20]. Geolocation [MSMS19]. Geospatial [BCJ20]. Global [FAC20]. Gradients [BCJ20]. Grind [JB18]. Group [LXB^{+}18, TH19, RFM^{+}18]. Groups [TH19]. Guest [RFM^{+}18]. Guide [MSMS19]. Gun [MADK22].

Happiness [WBUAV20]. Harmful [BYSB21]. Help [WFL^{+}19]. Helping [DMKH23]. HICSS [DK20]. Hidden [TH19]. Hospitality [RCRCOC^{+}19]. Human [GLL^{+}21, MYM23]. Human-centered [MYM23]. Hybrid [GP18]. Hypertargeting [Bay18].

Idea [SDCC19]. Identification [BYSB21]. Identifying [MWY^{+}18]. Identities [PLK^{+}21]. Identity [MZ21, YSBL18]. Immigration [EC22]. Impact [PLK^{+}21, TH19]. Impacts [FAC20]. Implicit [KZB^{+}20]. Incentives [FS19]. Inclusion [FAC20]. Increases [JCS^{+}21]. Independent [FBJF20]. Influence [EC22, MGS20]. Information [CRO^{+}19]. Informed [BYSB21]. Infrastructural [JB18]. Instances [SWK^{+}19]. Integration [SDCC19]. Interaction [WSH19]. Interactions [CSKD21]. Internet [FBJF20]. Internet-of-Things [FBJF20]. Interoperability [JB18]. Intersection [JB18]. Intrinsic [YSBL18]. Introduction [Cro18, DK20, LKTSA20]. Investigation [GLL^{+}21, KMZ^{+}21]. Investment [JM20]. Issue [CDK18, DK20, LKTSA20, RFM^{+}18]. Iterative [HMP^{+}21].

Language [BYSB21, FAC20]. Latent [MYM23]. Launch [WBUAV20]. Lean [ZSF22]. Learning [SRCGP18]. Lens [WBUAV20, ZWSS23]. Librarian [CDK18]. life [DMKH23]. Likert [AZ20]. Liners [KGK^{+}20]. Linkage [YSBL18]. Literature [BYSB21]. Living [FBJF20]. Locating [PLK^{+}21]. Location [PLK^{+}21]. Location-Based [PLK^{+}21]. Looking [SRCGP18].

Machine [WFL^{+}19]. Macrotasks [SL18]. Makes [ZYM^{+}21]. Management [GB20, SDCC19]. Marginality [EC22]. matter [YNMA20]. Me [WFL^{+}19]. Measuring [MGS20, PBL^{+}19]. Mechanism [SAM23]. Media [ARZ^{+}22, BCJ20, Bay18, EC22, FAC20, GDGM18, KMZ^{+}21, MADK22, MBWM21, RMG20, WGB20, YSBL18]. Meditation [JCS^{+}21]. Members [CSKD21]. Meme [FKW23]. Memory [PKK20]. Methods [JM20]. MFPR [SLZ^{+}18]. Mi [RCRCOC^{+}19]. Microblog [TPR^{+}18]. Microtask [FS19]. Mining [ARN20]. Mixed [JM20]. Mixed-Methods [JM20]. Mobile [ARN20, FEH20]. Modeling [MADK22, YSBL18]. Modifiers [RMG20]. Monetary [FS19]. Motivation [PBL^{+}19]. Motivations [PBL^{+}19]. Movements [MBWM21]. Multi [DCDCM19, LXB^{+}18]. Multi-Criteria [LXB^{+}18]. Multi-Tenant [DCDCM19]. Multidimensional [PBL^{+}19]. Multiplayer [KPKL23]. Multiple [SLZ^{+}18].

Narratives [PKK20]. Negotiating

[LKTSA20]. Network [CSKD21, KMV^{+}19]. Networking [SDCC19]. Networks [CRO^{+}19, LBZZ20, MGS20, PLK^{+}21]. Newcomer [TFLK18]. News [MWY^{+}18, PMY21]. No [KGK^{+}20]. Non [SL18]. Non-Decomposable [SL18].

Obtaining [KZB^{+}20]. Offensive [BYSB21]. Offline [KPKL23]. One [KGK^{+}20, ZYM^{+}21]. Online [AZ20, DMKH23, JM20, KPKL23, KGTA18, LRSW23, MGS20, MWY^{+}18, SDCC19, SdMAP20, SL18, TFLK18, TH19, WGB20, WSH19, ZWSS23]. Open [HR23]. opted [EC22]. Other [DMKH23]. Outcomes [DMKH23]. Overhauls [FEH20]. Overlapping [LBZZ20].

Paid [FS19]. Pandemic [LRSW23]. Paradigm [WSH19]. Participant [AZ20]. Participation [ARZ^{+}22, GP18]. Partisan [PMY21]. Patient [ARN20]. Patterns [SdMAP20]. Payment [JWZW18]. Payment-Driven [JWZW18]. Peer [ALL^{+}20, PKK20]. Peer-Produced [PKK20]. Peer-Production [ALL^{+}20]. Perception [FEH20, SRCGP18, TC20]. Perceptions [AZ20]. Performance [ZSF22]. Peripheral [GP18]. Permeability [JB18]. Personal [CG20]. Personalized [SLZ^{+}18, ZYM^{+}21]. Phishing [GLL^{+}21]. Ping [ZSF22]. Points [LXB^{+}18]. Political [Bay18, MADK22]. Politician [CDK18]. Politicians [SdMAP20]. Politics [SdMAP20]. Popular [KZB^{+}20]. Portal [ARN20]. Portrayal [KMZ^{+}21]. Postmortem [GB20]. Posts [ZYM^{+}21]. Powered [DCDCM19]. Practical [MSMS19]. Practice [KGTA18]. Practices [RCRCOC^{+}19]. Practitioners [KGTA18]. Predict [MGS20]. Presence [JCS^{+}21]. Presentations [PLK^{+}21]. Problems [HR23]. Process [HMP^{+}21, PMY21]. Processes [JM20]. Produced [PKK20]. Production [ALL^{+}20]. Proficiency [FAC20]. Profile [GB20]. Profiles [TH19].

Program [TPR^{+}18]. Propagation [CRO^{+}19]. Protection [KMV^{+}19]. Psychology [BYSB21]. Psychometrics [Bay18]. Public [TR18].

Q&A [JWZW18, SWK^{+}19]. Qualitative [AZ20]. Quantifying [GDGM18]. Question [SWK^{+}19]. Question-and-Answer [SWK^{+}19]. Quit [DMKH23].

r [DMKH23]. r/StopSmoking [DMKH23]. Ranking [SLZ^{+}18]. Rawlsian [Bay18]. Reactions [WBUAV20]. Reader [PMY21]. Real [DMKH23]. Real-life [DMKH23]. Reality [HMP^{+}21, JCS^{+}21]. Recommendation [SLZ^{+}18]. Recommendations [KGK^{+}20]. Reddit [KGTA18]. Regulation [KMV^{+}19]. Related [GLL^{+}21]. Relationship [KPKL23]. Representation [PMY21]. Requests [FLZ^{+}20]. Research [HMP^{+}21, KMV^{+}19]. Researchers [SAM23]. Respiratory [JCS^{+}21]. Response [TFLK18]. Responses [BCJ20]. Richness [FAC20]. Roles [KGTA18]. Routines [ALL^{+}20].

Satisfaction [FAC20]. Scale [PBL^{+}19]. Scales [AZ20]. Scheduling [DCDCM19]. Science [HMP^{+}21, SRP^{+}18, WFL^{+}19]. Self [LRSW23, MZ21, PLK^{+}21]. Self-Declared [MZ21]. Self-Disclosure [LRSW23]. Sensemaking [ZWSS23]. Sentiment [KZB^{+}20]. Sequences [ALL^{+}20]. Sequencing [SL18]. Shared [JCS^{+}21]. Sharing [SAM23]. Shipping [JB18]. Sites [PMY21]. Size [TH19]. Skin [RMG20]. Smoking [DMKH23]. Snapchat [FEH20]. Sociability [HMP^{+}21]. Social [ARZ^{+}22, BCJ20, Bay18, CRO^{+}19, Cro18, EC22, FBJF20, GDGM18, JCS^{+}21, KMZ^{+}21, KMV^{+}19, KGTA18, LRSW23, MGS20, MADK22, MBWM21, PLK^{+}21, PMY21, RMG20, SDCC19, WGB20, WSH19, YSBL18]. Socialization [TFLK18]. Socio [LKTSA20, TR18]. Socio-Contextual [TR18]. Socio-Technical [LKTSA20]. Sociopolitical [PKK20]. Solution [FBJF20]. South [SRCGP18]. Southeast [AC21]. Special [CDK18, DK20, LKTSA20, RFM^{+}18]. Spotlight [ZWSS23]. Stack [SWK^{+}19]. StopSmoking [DMKH23]. Streamers [AC21]. Study [AZ20, CG20, KGTA18, SDCC19, SWK^{+}19, TC20, ZSF22]. Subject [YNMA20]. Subject-matter [YNMA20]. Support [GP18, LRSW23]. Supporting [LXB^{+}18]. Survey [BYSB21]. Surveys [AZ20]. Susceptibility [GLL^{+}21]. Synchrony [JCS^{+}21]. Systematic [HR23]. Systems [CG20, DCDCM19, JWZW18, LKTSA20].

Talk [SdMAP20]. Targeted [JWZW18]. Task [FAC20]. Team [ZSF22]. Teams [FAC20]. Technical [BYSB21, GLL^{+}21, LKTSA20]. Temporal [ALL^{+}20, BCJ20]. Temporality [PLK^{+}21]. Tenant [DCDCM19]. Tendencies [KPKL23]. Terms [FEH20]. Terrorism [BCJ20]. Testing [MYM23]. Them [FLZ^{+}20]. Theory [ARZ^{+}22]. Things [FBJF20]. Three [HMP^{+}21]. Time [MGS20, PLK^{+}21]. Time-Constrained [MGS20]. Tone [RMG20]. Topic [WGB20]. Topical [MZ21]. Toxic [KPKL23]. Toxicity [MYM23]. Tracking [CG20]. Transactions [Cro18]. Transferability [MYM23]. Trust [GB20, JM20, LKTSA20, ZYM^{+}21]. Truth [LKTSA20]. Tweets [ZWSS23]. Twitch.tv [KZB^{+}20]. Twitter [CSKD21, FLZ^{+}20, KMZ^{+}21, MSMS19, TR18].

U.S. [CSKD21]. Uncertainty [CRO^{+}19]. Uncertainty-based [CRO^{+}19]. Understanding [AZ20, DMKH23, JM20, KGTA18]. Unpacking [MYM23]. Urban [SRCGP18].

Usage [RMG20, WGB20]. Use [ZSF22]. User [ARN20, DMKH23, FEH20, GP18, KGTA18, LRSW23, MSMS19, YSBL18]. User-generated [ARN20]. Users [FLZ^{+}20, WBUAV20]. Using [WBUAV20]. UX [CDK18].

Variation [RMG20]. Velocity [JB18].

Versions [HMP^{+}21]. Via

[JCS^{+}21, MADK22, WSH19]. Videos [WSH19]. Viewer [KZB^{+}20]. Viral [ARZ^{+}22]. Virtual [FAC20, HMP^{+}21, JCS^{+}21, SRP^{+}18]. Visual [PKK20]. Visualizing [LXB^{+}18]. Volatile [KGTA18].

Walk [CDK18]. Web [BYSB21]. Websites [SWK^{+}19]. Weighted [LBZZ20]. Where [FLZ^{+}20]. Wikipedia [ALL^{+}20, MZ21, YNMA20]. Wisely [ZSF22]. within [ZWSS23]. Women [AC21]. Work [ALL^{+}20, TPR^{+}18].

Years [RFM^{+}18]. Young [ARZ^{+}22]. YouTube [KMZ^{+}21].

References

Alvarez:2021:CCE

[AC21] Katrina Paola B. Alvarez and Vivian Hsueh Hua Chen. Community and capital: Experiences of women game streamers in Southeast Asia. ACM Transactions on Social Computing (TSC), 4(3): 12:1–12:22, September 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                                     dl.acm.org/doi/10.1145/
                                                     3481888.

Arazy:2020:ERP

[ALL^{+}20] Ofer Arazy, Aron Lindberg, Shakked Lev, Kexian Wu, and Alex Yarovoy. Emergent routines in peer-production: Examining the temporal evolution of Wikipedia’s work sequences. ACM Transactions on Social Computing (TSC), 3(1):5:1–5:24, February 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3366711.

Al-Ramahi:2020:MUG

[ARN20] Mohammad Al-Ramahi and Cherie Noteboom. Mining user-generated content of mobile patient portal: Dimensions of user experience. ACM Transactions on Social Computing (TSC), 3(3):15:1– 15:24, August 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
10.1145/3394831.

Abraham:2022:ABC

[ARZ^{+}22] Jaclyn Abraham, Rebecca Roth, Heidi Zinzow, Kapil Chalil Madathil, and Pamela Wisniewski. Applying behavioral contagion theory to examining young adults’ participation in viral social media challenges. ACM Transactions on Social Computing (TSC), 5(1–4):3:1–3:??, December 2022. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3538383.

Alismail:2020:EUP

[AZ20] Sarah Alismail and Hengwei Zhang. Exploring and understanding participants’ perceptions of facial emoji Likert scales in online surveys: a qualitative study. ACM Transactions on Social Computing (TSC), 3(2): 12:1–12:12, April 2020. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3382505.

Bay:2018:SME

[Bay18] Morten Bay. Social media ethics: a Rawlsian approach to hypertargeting and psychometrics in political and commercial campaigns. ACM Transactions on Social Computing (TSC), 1(4): 16:1–16:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3281450.

Baucum:2020:TGG

[BCJ20] Matthew Baucum, Jinshu Cui, and Richard S. John. Temporal and geospatial gradients of fear and anger in social media responses to terrorism. ACM Transactions on Social Computing (TSC), 2 (4):16:1–16:16, January 2020.

CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3363565.

Balayn:2021:AIH

[BYSB21] Agathe Balayn, Jie Yang, Zoltan Szlavik, and Alessandro Bozzon. Automatic identification of harmful, aggressive, abusive, and offensive language on the Web: a survey of technical biases informed by psychology literature. ACM Transactions on Social Computing (TSC), 4(3):11:1–11:56, September 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3479158.

Crowston:2018:LPU

[CDK18] Kevin Crowston, Xuefei (Nancy)

Deng, and Yoram M. Kalman. A librarian, a politician, a UX expert, and a cyberbully walk into a special issue. ACM Transactions on Social Computing (TSC), 1(4): 14:1–14:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3293613.

Cifor:2020:GDD

[CG20] Marika Cifor and Patricia Garcia. Gendered by design: a duoethnographic study of personal fitness tracking systems. ACM Transactions on Social Computing (TSC), 2 (4):15:1–15:22, January 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3364685.

Crowston:2018:IAT

[Cro18]  Kevin Crowston.  Intro-

duction to ACM Transactions on Social Computing. ACM Transactions on Social Computing (TSC), 1(1):1:1– 1:??, February 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3181713.

Cho:2019:UBF

[CRO^{+}19] Jin-Hee Cho, Scott Rager, John O’Donovan, Sibel Adali, and Benjamin D. Horne. Uncertainty-based false information propagation in social networks. ACM Transactions on Social Computing (TSC), 2(2):5:1–5:??, October 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3311091.

Chamberlain:2021:NAT

[CSKD21] Joshua M. Chamberlain, Francesca Spezzano, Jaclyn J. Kettler, and Bogdan Dit. A network analysis of Twitter interactions by members of the U.S. Congress.

ACM Transactions on Social Computing (TSC), 4(1):1:1– 1:22, April 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
10.1145/3439827.

Difallah:2019:DAF

[DCDCM19] Djellel Difallah, Alessandro Checco, Gianluca Demartini, and Philippe Cudré-Mauroux. Deadline-aware fair scheduling for multitenant crowd-powered systems. ACM Transactions on Social Computing (TSC), 2 (1):3:1–3:??, February 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3301003.

Deng:2020:ISI

[DK20] Xuefei (Nancy) Deng and Yoram M. Kalman. Introduction to the special issue on HICSS 2019. ACM Transactions on Social Computing (TSC), 2(4):13:1–13:2, January 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3370666.

DeSanto:2023:HEO

[DMKH23] Alessio De Santo, Arielle Moro, Bruno Kocher, and Adrian Holzer. Helping each other quit online: Understanding user engagement and real-life outcomes of the r/StopSmoking digital smoking cessation community. ACM Transactions on Social Computing (TSC), 6(1– 2):1:1–1:??, June 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3564745.

Emes:2022:COM

[EC22] Claire Stravato Emes and Arul Chib. Co-opted marginality in a controlled media environment: The influence of social media affordances on the immigration discourse. ACM Transactions on Social Computing (TSC), 5(1–4):1:1– 1:??, December 2022. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3532103.

Fleischmann:2020:LPM

[FAC20] A. Carolin Fleischmann, Jolanta Aritz, and Peter Cardon. Language proficiency and media richness in global virtual teams: Impacts on satisfaction, inclusion, and task accomplishment. ACM Transactions on Social Computing (TSC), 2 (4):17:1–17:18, January 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3363564.

Forkan:2020:ITS

[FBJF20] Abdur Rahim Mohammad Forkan, Philip Branch, Prem Prakash Jayaraman, and Andre Ferretto. An Internet-of-Things solution to assist independent living and social connectedness in elderly. ACM Transactions on Social Computing (TSC), 2(4):14:1–14:24, January 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3363563.

Franzmann:2020:HMA

[FEH20] Daniel Franzmann, Arvid Eichner, and Roland Holten. How mobile app design overhauls can be disastrous in terms of user perception: The case of snapchat. ACM Transactions on Social Computing (TSC), 3(4):18:1–18:21, October 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3409585.

Ford:2023:CDM

[FKW23] Trenton W. Ford, Rachel Krohn, and Tim Weninger. Competition dynamics in the meme ecosystem. ACM Transactions on Social Computing (TSC), 6(3–4):6:1– 6:??, December 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3596213.

Feng:2020:NER

[FLZ^{+}20] Yunhe Feng, Zheng Lu, Wenjun Zhou, Zhibo Wang, and Qing Cao. New emoji requests from Twitter users: When, where, why, and what we can do about them. ACM Transactions on Social Computing (TSC), 3(2):8:1–8:25, April 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3370750.

Feyisetan:2019:BMI

[FS19] Oluwaseyi Feyisetan and Elena Simperl. Beyond monetary incentives: Experiments in paid microtask contests. ACM Transactions on Social Computing (TSC), 2(2):6:1– 6:??, October 2019. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3321700.

Gach:2020:ETP

[GB20] Katie Z. Gach and Jed R. Brubaker. Experiences of trust in postmortem profile management. ACM Transactions on Social Computing (TSC), 3(1):2:1–2:26, February 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3365525.

Garimella:2018:QCS

[GDGM18] Kiran Garimella, Gianmarco De Francisci Morales, Aristides Gionis, and Michael Mathioudakis. Quantifying controversy on social media. ACM Transactions on Social Computing (TSC), 1(1):3:1– 3:??, February 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3140565.

Greitzer:2021:EIT

[GLL^{+}21] Frank L. Greitzer, Wanru Li, Kathryn B. Laskey, James Lee, and Justin Purl. Experimental investigation of technical and human factors related to phishing susceptibility. ACM Transactions on Social Computing (TSC), 4(2):8:1–8:48, June 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3461672.

Gasson:2018:PAS

[GP18] Susan Gasson and Michelle Purcelle. A participation architecture to support user peripheral participation in a hybrid FOSS community. ACM Transactions on Social Computing (TSC), 1(4): 14:1–14:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3290837.

Holopainen:2021:SVR

[HMP^{+}21] Jani Holopainen, Osmo Mattila, Petri Parvinen, Essi Pöyry, and Tuure Tuunanen. Sociability in virtual reality: Evaluations of three iterative application versions along a design science research process. ACM Transactions on Social Computing (TSC), 4(1):4:1–4:21, April 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450269.

Heltweg:2023:SAP

[HR23] Philip Heltweg and Dirk Riehle. A systematic analysis of problems in open collaborative data engineering. ACM Transactions on Social Computing (TSC), 6(3–4):8:1– 8:??, December 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3629040.

Jabbar:2018:PIV

[JB18] Karim Jabbar and Pernille Bjørn. Permeability, interoperability, and velocity: Entangled dimensions of infrastructural grind at the intersection of blockchain and shipping. ACM Transactions on Social Computing (TSC), 1(3): 10:1–10:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3288800.

Jarvela:2021:AVR

[JCS^{+}21] Simo Järvelä, Benjamin Cowley, Mikko Salminen, Giulio Jacucci, Juho Hamari, and Niklas Ravaja. Augmented virtual reality meditation: Shared dyadic biofeedback increases social presence via respiratory synchrony. ACM Transactions on Social Computing (TSC), 4(2):6:1–6:19, June 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
10.1145/3449358.

Jones:2020:MMA

[JM20] Helen S. Jones and Wendy Moncur. A mixed-methods approach to understanding funder trust and due diligence processes in online crowdfunding investment. ACM Transactions on Social Computing (TSC), 3(1):3:1–3:29, February 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3373148.

Jan:2018:APD

[JWZW18] Steve T. K. Jan, Chun Wang, Qing Zhang, and Gang Wang. Analyzing paymentdriven targeted Q&A systems. ACM Transactions on Social Computing (TSC), 1(3):12:1–12:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/citation.cfm?
                                        id=3281449.

Kim:2020:NMO

[KGK^{+}20] Joongyum Kim, Taesik Gong, Bogoan Kim, Jaeyeon Park, Woojeong Kim, Evey Huang, Kyungsik Han, Juho Kim, Jeonggil Ko, and Sung-Ju Lee. No more one liners: Bringing context into emoji recommendations. ACM Transactions on Social Computing (TSC), 3(2):9:1–9:25, April 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3373146.

Kou:2018:USR

[KGTA18] Yubo Kou, Colin M. Gray, Austin L. Toombs, and Robin S. Adams. Understanding social roles in an online community of volatile practice: A study of user experience practitioners on Reddit. ACM Transactions on Social Computing (TSC), 1(4): 17:1–17:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3283827.

Kotsios:2019:ACG

[KMV^{+}19] Andreas Kotsios, Matteo Magnani, Davide Vega, Luca Rossi, and Irina Shklovski. An analysis of the consequences of the General Data Protection Regulation on social network research. ACM Transactions on Social Computing (TSC), 2(3): 12:1–12:22, December 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365524.

Khasawneh:2021:IPS

[KMZ^{+}21] Amro Khasawneh, Kapil Chalil Madathil, Heidi Zinzow, Pamela Wisniewski, Amal Ponathil, Hunter Rogers, Sruthy Agnisarman, Rebecca Roth, and Meera Narasimhan. An investigation of the portrayal of social media challenges on YouTube and Twitter. ACM Transactions on Social Computing (TSC), 4(1):2:1–2:23, April 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444961.

Kordyaka:2023:ERB

[KPKL23] Bastian Kordyaka, Solip Park, Jeanine Krath, and Samuli Laato. Exploring the relationship between offline cultural environments and toxic behavior tendencies in multiplayer online games. ACM Transactions on Social Computing (TSC), 6(1– 2):3:1–3:??, June 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3580346.

Kobs:2020:ECO

[KZB^{+}20] Konstantin Kobs, Albin Zehe, Armin Bernstetter, Julian Chibane, Jan Pfister, Julian Tritscher, and Andreas Hotho. Emotecontrolled: Obtaining implicit viewer feedback through emote-based sentiment analysis on comments of popular Twitch.tv channels. ACM Transactions on Social Computing (TSC), 3(2):7:1–7:34, April 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3365523.

Lyu:2020:FFO

[LBZZ20] Tianshu Lyu, Lidong Bing, Zhao Zhang, and Yan Zhang. FOX: Fast overlapping community detection algorithm in big weighted networks. ACM Transactions on Social Computing (TSC), 3(3):16:1– 16:23, August 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3404970.

Luther:2020:ISI

[LKTSA20] Kurt Luther, Andrea Kavanaugh, Jacob Thebault-Spieker, and Judd Antin. Introduction to the special issue on negotiating truth and trust in socio-technical systems. ACM Transactions on Social Computing (TSC), 3 (1):1:1, February 2020. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3378677.

Lee:2023:OSD

[LRSW23] Jooyoung Lee, Sarah Rajtmajer, Eesha Srivatsavaya, and Shomir Wilson. Online selfdisclosure, social support, and user engagement during the COVID-19 pandemic. ACM Transactions on Social Computing (TSC), 6(3–4):7:1– 7:??, December 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3617654.

Liu:2018:CSM

[LXB^{+}18] Weichen Liu, Sijia Xiao, Jacob T. Browne, Ming Yang, and Steven P. Dow. ConsensUs: Supporting multicriteria group decisions by visualizing points of disagreement. ACM Transactions on Social Computing (TSC), 1 (1):4:1–4:??, February 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3159649.

Mejova:2022:MPA

[MADK22] Yelena Mejova, Jisun An, Gianmarco De Francisci Morales, and Haewoon Kwak. Modeling political activism around gun debate via social media. ACM Transactions on Social Computing (TSC), 5(1–4):2:1–2:??, December 2022. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3532102.

Mirbabaie:2021:DCA

[MBWM21] Milad Mirbabaie, Felix Brünker, Magdalena Wischnewski, and Judith Meinert. The development of connective action during social movements on social media. ACM Transactions on Social Computing (TSC), 4(1):3:1–3:21, April 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3446981.

Marin:2020:MTC

[MGS20] Ericsson Marin, Ruocheng Guo, and Paulo Shakarian. Measuring time-constrained influence to predict adoption in online social networks. ACM Transactions on Social Computing (TSC), 3(3):13:1– 13:26, August 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3372785.

Mourad:2019:PGE

[MSMS19] Ahmed Mourad, Falk Scholer, Walid Magdy, and Mark Sanderson. A practical guide for the effective evaluation of Twitter user geolocation. ACM Transactions on Social Computing (TSC), 2 (3):9:1–9:23, December 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3352572.

Morstatter:2018:IFB

[MWY^{+}18] Fred Morstatter, Liang Wu, Uraz Yavanoglu, Stephen R. Corman, and Huan Liu. Identifying framing bias in online news. ACM Transactions on Social Computing (TSC), 1 (2):5:1–5:??, June 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3204948.

Muralikumar:2023:HCE

[MYM23] Meena Devii Muralikumar, Yun Shan Yang, and David W. McDonald. A humancentered evaluation of a toxicity detection API: Testing transferability and unpacking latent attributes. ACM Transactions on Social Computing (TSC), 6(1–2):4:1– 4:??, June 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3582568.

Mcdonald:2021:ABS

[MZ21] David W. Mcdonald and Mark Zachry. On the alignment between self-declared gender identity and topical content from Wikipedia.

ACM Transactions on Social Computing (TSC), 4(2):7:1– 7:69, June 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450753.

Posch:2019:MMC

[PBL^{+}19] Lisa Posch, Arnim Bleier, Clemens M. Lechner, Daniel Danner, Fabian Flöck, and Markus Strohmaier. Measuring motivations of crowdworkers: The multidimensional crowdworker motivation scale. ACM Transactions on Social Computing (TSC), 2(2):8:1–8:??, October 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3335081.

Porter:2020:VNC

[PKK20] Emily Porter, P. M. Krafft, and Brian Keegan. Visual narratives and collective memory across peer-produced accounts of contested sociopolitical events. ACM Transactions on Social Computing (TSC), 3(1):4:1–4:20, February 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3373147.

Papangelis:2021:LIT

[PLK^{+}21] Konstantinos Papangelis, Ioanna Lykourentzou, Vassilis-Javed Khan, Alan Chamberlain, Ting Cao, Michael Saker, and Nicolas Lalone. Locating identities in time: an examination of the formation and impact of temporality on presentations of the self through location-based social networks. ACM Transactions on Social Computing (TSC), 4(3):10:1–10:23, September 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3473043.

Purao:2021:DBS

[PMY21] Sandeep Purao, David M.

Murungi, and David Yates. Deliberative breakdowns in the social representation process: Evidence from reader comments in partisan news sites. ACM Transactions on Social Computing (TSC), 4 (2):5:1–5:35, June 2021. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

dl.acm.org/doi/10.1145/
3450143.

Ruiz-Correa:2019:MCC

[RCRCOC^{+}19] Salvador Ruiz-Correa, Itzia Ruiz-Correa, Carlo Olmos-Carrillo, Fatima Alba Rendón-Huerta, Beatriz Ramirez-Salazar, Laurent Son Nguyen, and Daniel Gatica-Perez. Mi casa es su casa? Examining Airbnb hospitality exchange practices in a developing economy. ACM Transactions on Social Computing (TSC), 2(1):2:1–2:??, February 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/citation.cfm?
                                        id=3299817.

Robert:2018:GSI

[RFM^{+}18] Lionel P. Robert, Jr., Andrea Forte, Claudia Müller, Michael Prilla, and Adriana S. Vivacqua. GROUP 2018 special issue guest editorial: Another 25 years of GROUP. ACM Transactions on Social Computing (TSC), 1(3):9:1–9:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3290870.

Robertson:2020:EST

[RMG20] Alexander Robertson, Walid Magdy, and Sharon Goldwater. Emoji skin tone modifiers: Analyzing variation in usage on social media. ACM Transactions on Social Computing (TSC), 3(2): 11:1–11:25, April 2020. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3377479.

Seneviratne:2023:ABB

[SAM23] Oshani Seneviratne, Kacy Adams, and Deborah L. McGuinness. Accountable bench-to-bedside datasharing mechanism for researchers. ACM Transactions on Social Computing (TSC), 6(3–4):5:1–5:??, December 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609486.

Saldivar:2019:OIM

[SDCC19] Jorge Saldivar, Florian Daniel, Luca Cernuzzi, and Fabio Casati. Online idea management for civic engagement: A study on the benefits of integration with social networking. ACM Transactions on Social Computing (TSC), 2 (1):1:1–1:??, February 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3284982.

SantosDeOliveira:2020:DPT

[SdMAP20] Lucas Santos De Oliveira, Pedro O. S. Vaz de Melo, Marcelo S. Amaral, and José Antônio G. Pinho. Do politicians talk about politics? Assessing online communication patterns of Brazilian politicians. ACM Transactions on Social Computing (TSC), 3(4):19:1–19:28, October 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3412326.

Schmitz:2018:OSN

[SL18] Heinz Schmitz and Ioanna Lykourentzou. Online sequencing of non-decomposable macrotasks in expert crowdsourcing. ACM Transactions on Social Computing (TSC), 1(1):1:1–1:??, February 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3140459.

Shi:2018:MPR

[SLZ^{+}18] Chuan Shi, Jian Liu, Yiding Zhang, Binbin Hu, Shenghua Liu, and Philip S. Yu. MFPR: A personalized ranking recommendation with multiple feedback. ACM Transactions on Social Computing (TSC), 1(2):7:1–7:??, June 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3216368.

Santani:2018:LSL

[SRCGP18] Darshan Santani, Salvador Ruiz-Correa, and Daniel Gatica-Perez. Looking south: Learning urban perception in developing cities. ACM Transactions on Social Computing (TSC), 1(3):13:1– 13:??, December 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3224182.

Simperl:2018:VCS

[SRP^{+}18] Elena Simperl, Neal Reeves, Chris Phethean, Todd Lynes, and Ramine Tinati. Is virtual citizen science a game? ACM Transactions on Social Computing (TSC), 1(2): 6:1–6:??, June 2018. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3209960.

Santos:2019:AAQ

[SWK^{+}19] Tiago Santos, Simon Walk, Roman Kern, Markus Strohmaier, and Denis Helic. Activity archetypes in question-and-answer (Q&A) Websites — a study of 50 Stack Exchange instances. ACM Transactions on Social Computing (TSC), 2(1):4:1–4:??, February 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3301612.

Talukder:2020:SFA

[TC20] Sajedul Talukder and Bogdan Carbunar. A study of friend abuse perception in Facebook. ACM Transactions on Social Computing (TSC), 3(4):17:1–17:34, October 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3408040.

Tausczik:2018:ECS

[TFLK18] Yla Tausczik, Rosta Farzan, John Levine, and Robert Kraut. Effects of collective socialization on newcomers’ response to feedback in online communities. ACM Transactions on Social Computing (TSC), 1(2):8:1–8:??, June 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/citation.cfm?
                                        id=3191834.

Tausczik:2019:IGS

[TH19] Yla Tausczik and Xiaoyun Huang. The impact of group size on the discovery of hidden profiles in online discussion groups. ACM Transactions on Social Computing (TSC), 2(3):10:1–10:??, November 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/citation.cfm?
                                        id=3359758.

Tolmie:2018:MAP

[TPR^{+}18] Peter Tolmie, Rob Procter, Mark Rouncefield, Maria Liakata, and Arkaitz Zubiaga. Microblog analysis as a program of work. ACM Transactions on Social Computing (TSC), 1(1):2:1–2:??, February 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

                                        dl.acm.org/citation.cfm?
                                        id=3162956.

Tahmasbi:2018:SCA

[TR18] Nargess Tahmasbi and Elham Rastegari. A sociocontextual approach in automated detection of public cyberbullying on Twitter. ACM Transactions on Social Computing (TSC), 1(4): 15:1–15:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3290838.

Wisniewski:2020:HFU

[WBUAV20] Pamela Wisniewski, Karla Badillo-Urquiola, Zahra Ashtorab, and Jessica Vitak. Happiness and fear: Using emotions as a lens to disentangle how users felt about the launch of Facebook reactions. ACM Transactions on Social Computing (TSC), 3(4):20:1–20:25, October 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414825.

Wright:2019:HMH

[WFL^{+}19] Darryl E. Wright, Lucy Fortson, Chris Lintott, Michael Laraia, and Mike Walmsley. Help me to help you: Machine augmented citizen science. ACM Transactions on Social Computing (TSC), 2(3): 11:1–11:??, November 2019. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3362741.

Wood:2020:CTU

[WGB20] Ian D. Wood, John Glover, and Paul Buitelaar. Community topic usage in online social media. ACM Transactions on Social Computing (TSC), 3(3):14:1–14:21, August 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3377870.

Wu:2019:DNP

[WSH19] Qunfang Wu, Yisi Sang, and Yun Huang. Danmaku: A new paradigm of social interaction via online videos. ACM Transactions on Social Computing (TSC), 2(2):7:1– 7:??, October 2019. CO-DEN ???? ISSN 2469-7818 (print), 2469-7826 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3329485.

Yarovoy:2020:ACS

[YNMA20] Alex Yarovoy, Yiftach Nagar, Einat Minkov, and Ofer Arazy. Assessing the contribution of subject-matter experts to Wikipedia. ACM Transactions on Social Computing (TSC), 3(4):21:1– 21:36, October 2020. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3416853.

Yu:2018:MUI

[YSBL18] Xianqi Yu, Yuqing Sun, Elisa Bertino, and Xin Li. Modeling user intrinsic characteristic on social media for identity linkage. ACM Transactions on Social Computing (TSC), 1(3):11:1–11:??, December 2018. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL https://

dl.acm.org/citation.cfm?
id=3267442.

Zheng:2022:UPW

[ZSF22] Keyang Zheng, Ben Stein, and Rosta Farzan. Use ping wisely: a study of team communication and performance under lean affordance. ACM Transactions on Social Computing (TSC), 5(1–4):4:1– 4:??, December 2022. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3557022.

Zhou:2023:STL

[ZWSS23] Kaitlyn Zhou, Tom Wilson, Kate Starbird, and Emma S. Spiro. Spotlight tweets: a lens for exploring attention dynamics within online sensemaking during crisis events. ACM Transactions on Social Computing (TSC), 6(1– 2):2:1–2:??, June 2023. ISSN 2469-7818 (print), 2469-7826 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577213.

Zhou:2021:BPT

[ZYM^{+}21] Michelle X. Zhou, Huahai Yang, Gloria Mark, Mengdie Hu, Jalal Mahumd, and Aditya Pal. Building personalized trust: Discovering what makes one trust and act on Facebook posts. ACM Transactions on Social Computing (TSC), 4(3):9:1–9:28, September 2021. CODEN ???? ISSN 2469-7818 (print), 2469-7826 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3468977.